OpenStudy (anonymous):

Had to solve for v |5v-15|=10 I got two answers. (1. -5,5) (2. 5,3) Can someone tell me which one is right?

OpenStudy (anonymous):

when finding absolute value, you need to form it into 2 equations

OpenStudy (anonymous):

One where you set the answer as a negative, and another where you set it as a positive

OpenStudy (anonymous):

Did you do that?

OpenStudy (anonymous):

yes but I did -5v-15=10\5v-15=10 and the other 5v-15=-10\5v-15=10 because I was shown these two ways but I am not sure which is right in this case.

OpenStudy (anonymous):